# Towards improved property prediction of two-dimensional (2D) materials using many-body Quantum Monte Carlo methods

### v1

@@ II Theory
Although various real-space DMC codes have been developed over the last few decades and applied to 2D materials (e.g., CASINO [142], PyQMC [143], QWalk [144]), most of the work in this review article utilizes the QMCPACK code[145, 146]. The QMCPACK code has been routinely updated and maintained to track the latest algorithmic and hardware developments and utilize computational resources more efficiently (https://github.com/QMCPACK/qmcpack). The QMCPACK code is equipped with the Nexus [147] workflow automation software, which helps create and monitor DFT/DMC workflows that cansignificantly reduce the user time by tracking computational dependencies. More detailed information about these theoretical frameworks and codes can be found in Refs.103, 63, 119, 120, 145, 146, 142, 143, 144, 148. For this review article, it is important to note that the error bars in each figure and the uncertainty noted in in concise notation (i.e., 5.3 0.1 can be expressed as 5.3(1)) refer to the standard error about the mean for each observable quantity calculated with QMC (i.e., energy, band gap, magnetic moment, lattice constant).
@@ III.1 Magnetic Properties
After the experimental synthesis of monolayer CrI3, which was measured to have a Curie temperature (Tc) of 45 K [16], significant interest in identifying and physically understanding similar magnetic 2D materials has increased. It has been experimentally reported that monolayer VSe2 is ferromagnetic on a vdW substrate [149] and that ferromagnetic order exists in the bilayer limit for Cr2Ge2Te6 [150]. Some other examples of low-dimensional magnetic systems include FePS3 [151], NiPS3 [152], MnPS3 [153], and Fe3GeTe2 [154]. Several computational works have also predicted magnetism in a variety of 2D systems [155, 156, 157, 158].
The Mermin-Wagner theorem [159] states that magnetic order in a 2D material cannot persist unless magnetic anisotropy (MA) is present. Moreover, it requires anisotropy that breaks continuous symmetries. Whether the anisotropy is perpendicular to the plane (easy-axis) or parallel to the plane (easy plane) determines the type of transition temperature. For easy-axis anisotropy, the system can be described by the Heisenberg model with a finite Curie temperature . For easy-plane anisotropy, there is no explicit transition between the ordered and unordered states, but instead a Kosterlitz-Thouless transition at finite temperature , where the system has quasi-long-range magnetic order below . To obtain accurate estimates for observables such as and , an accurate computation of the magnetic exchange and magnetic anisotropy energies is essential.Oftentimes, these results are extremely sensitive to the choice of density functional and Hubbard U parameter, resulting in inconclusive results with DFT. Additionally, these 2D magnetic materials can have a strong interdependence between structural parameters and magnetic properties, making it even more difficult to make accurate predictions. Due to this, using a method such as DMC to understand the electron correlation effects that drive magnetic ordering in 2D is necessary. These accurately computed ab initio magnetic parameters can then be used in conjunction with either classical Monte Carlo methods or other analytical models [160] to estimate observable quantities such as transition temperatures. Next, we will discuss few of our works on modeling magnetic 2D structures.
@@ III.1.1 Monolayer CrI3
One of the first monolayer magnets discovered,[16] monolayer CrI3 is a Mott insulator [161] that consists of chromium atoms octahedrally coordinated by six iodine atoms (see Figures 1 and 2). Because of the negative exchange interactions among them, the spins on the Cr atoms arrange themselves perpendicular to the plane of the material,[162, 163] resulting in a monolayer ferromagnet with a critical temperature of 45 K.[16, 164] Upon stacking, CrI3 heterostructures oscillate between exhibiting ferromagnetism for odd numbers of layers and antiferromagnetism for even numbers of layers.[165] Moreover, upon the application of pressure that decreases the interlayer distance, bilayer CrI3 transitions from an antiferromagnet to a ferromagnet.[166, 167]
@@
These phenomena suggest that electron-electron and electron-phonon interactions play a key role in determining CrI3’s magnetic and structural properties. Yet, while significant research has gone into characterizing and modifying the macroscopic properties of this material, far less research has focused on the atomistic determinants of those properties in a manner that accurately accounts for many-body interactions. As a result, only approximate explanations for the source of this material’s magnetism had been put forth before the DMC research described below was conducted. In particular, based on a rudimentary analysis of oxidation states, it was originally believed that each iodine atom in CrI3 would have an oxidation state of -1 and therefore that the chromium should have an oxidation state of +3. Given that neutral Cr has a [Ar]4s13d5 electron configuration, this suggests that Cr3+ should have three valence electrons remaining, giving rise to a magnetic moment of 3 . More accurate, yet still single reference DFT calculations corroborated this picture – on average – but often differed significantly from one another in detail.[168, 169, 170, 162, 171] Previous PBE+U (U = 2 eV) and HSE calculations, for example, predicted Cr’s magnetic moment to be as large as 3.3 ,[169, 171] while PBE[171] and LDA+U [162] calculations predicted moments as small as 3.1 and 3 , respectively – a 10 % difference among moments overall. Even larger differences could be observed among DFT predictions of the in-plane lattice parameter: PBE calculations predicted a CrI3 lattice parameter of 7.008 Å,[171] while LDA+U calculations predicted a lattice parameter of 6.686 Å,[162] a 5 % difference. These significant DFT discrepancies were further corroborated by our own DFT calculations, which not only demonstrated that the magnetic moments and lattice parameters predicted monotonically increased with the U employed in PBE+U calculations (suggesting that it would be difficult to determine a meaningful U value from DFT calculations alone), but also that the U that most closely matched experiments for one property significantly differed from those needed to match experimental predictions of other properties.[109] Thus, while these predictions provide general insights into CrI3’s magnetism, they lack the accuracy – and precision – needed to make definitive statements.
@@
To bring clarity to this picture, we thus performed DMC simulations on monolayer CrI3, one of the first performed on a magnetic monolayer material.[109] Given the known sensitivity of the electronic structure of 2D materials to strain[172, 173] and the lack of experimental results for the monolayer geometry, however, we first set out to determine a high-accuracy, many-body monolayer geometry (see Fig. 1). To do so, we leveraged a cutting-edge, many-body geometry optimization method termed the surrogate Hessian line-search method.[174, 175, 176] In this approach, a Hessian generated using DFT is used to determine conjugate directions along which to locate the minimum-energy structure. Differing from standard DFT-based conjugate gradient minimization, the surrogate Hessian line-search method then computes DMC energies in parallel along each of the conjugate directions and subsequently refines the conjugate directions until the search converges to the DMC energyminimum.[174, 175, 176] This method is particularly advantageous because it makes the most efficient use of both DFT and DMC data: instead of relying on more expensive DMC energy gradients, it uses relatively cheap DFT gradients and then employs DMC energy calculations to refine the search, thus providing a substantially more efficient way of determining fully many-body geometries. Applying this approach to CrI3, we obtained a ground state geometry with a lattice parameter of Å and a Cr-I bond distance of Å .[109] Interestingly, while we were conducting this research, some of the first STM-based measurements of the monolayer geometry were performed yielding a lattice constant of Å.[168] Such strong agreement ( error) highlights the accuracy of our structure and the value of the surrogate Hessian line-search approach, despite CrI3’s (and many monolayer materials’) shallow potential energy surface around its minimum.
@@
With this high-accuracy structure in hand, we then performed DMC calculations on the monolayer. To generate a high-quality trial wavefunction from which to construct our DMC nodal surface, we first performed PBE+U calculations. Unlike previous works, we were sure to determine our U in a variational fashion, by determining the U in our PBE+U trial wavefunctions that minimized our resultant DMC energies (see Fig. 1). This resulted in an optimized U value of 2 eV.[109] Employing this U in our Slater-Jastrow trial wavefunctions with one- and two-body Jastrow terms, we obtained high-accuracy DMC charge and spin densities. We finally obtained DMC estimates of the Cr magnetic moment by integrating the spin density from the center of the Cr atom to the distance from the nucleus at which the spin density changed sign (the zero-recrossing radius of the sign of the spin density). This yielded a DMC magnetic moment of 3.62 , a moment substantially larger than that previously predicted using single reference techniques. Importantly, PBE+U=2 eV calculations performed on the same optimized geometries yielded a similarly large Cr moment of 3.57 , demonstrating the strong coupling between the moment and the geometry,[173, 172] as also observed in previous DFT simulations and our subsequent spin-phonon and spin-lattice coupling calculations.[109] Although these moments seem unexpectedly large, they are counterbalanced by negative moments of -0.145 on the iodine atoms, leading to the anticipated averaged moment of 3 over the entire unit cell. These simulations have therefore shed a clarifying light on the atomic-scale origin of CrI3’s magnetism and produced a valuable workflow that can be readily extended to a wide range of other 2D materials.[177] Additional details of this work can be found in Ref. 109.
@@ III.1.2 Monolayer CrX3
To expand upon this seminal work, we created a workflow that combines DFT+U and DMC methods to calculate magnetic properties of 2D CrX3 (X = I, Br, Cl, F) with improved accuracy [110]. These structures were chosen as a case study since they have been extensively investigated with DFT [178], have been experimentally synthesized, and have a finite transition temperature [16, 178]. We can map our ab initio quantities to a 2D model spin Hamiltonian to obtain properties such as [160, 162]. The model Hamiltonian contains easy axis single ion anisotropy (), Heisenberg isotropic exchange (), and anisotropic exchange (),
@@
The sum over runs over the lattice of Cr atoms (magnetic atoms) and runs over ’s nearest-neighbor Cr atoms; only nearest-neighbors are needed because the magnetic moments are strongly localized on the Cr atoms. prefers off-plane easy axis, prefers FM interactions, and results in fully isotropic exchange. This model Hamiltonian can additionally contain terms for Dzyaloshinskii-Moriya interactions and Kitaev interactions, but for this work, we limit our Hamiltonian to containing , , and terms. Note that in the following sections refers to the Heisenberg exchange and not the J in the DFT+U+J formalism [179]. Two magnetic orientations can be constructed from the CrX3 unit cell (shown in Fig. 2b): the ferromagnetic (FM) (two spin up Cr atoms) and antiferromagnetic (AFM) (one spin up and one spin down Cr atoms) orientations. , , and can be extracted from noncollinear (spin-orbit) DFT calculations. Specifically, computing the energy differences between configurations that have an easy axis rotation of 90∘ and non-rotated configurations for FM and AFM magnetic orientations.
@@
These results can be systematically improved with DMC and a full schematic of this workflow is shown in Fig. 2(a). Firstly, the nodal surface of the trial wavefunction can be optimized by changing the U value (variationally determining which U value yields the lowest energy). Secondly, by performing DMC for the FM and AFM configurations, can be estimated. In our case, these DMC energies are collinear (without spin-orbit coupling). Although SOC has been implemented in DMC, the energy differences needed to compute magnetic anisotropy parameters are smaller in magnitude than the DMC uncertainty. For this reason, we used the optimal U value determined from the nodal surface optimization (in our case, it was U = 2 eV for CrI3 and CrBr3, similar to Ref. 109) to perform DFT+U calculations to obtain and . The starting geometry for our DMC calculations was obtained from the vdW-DF-optB88 [79] functional (vdW-DF-optB88 produces a geometry for CrI3 in near-perfect agreement with experiment [168] and Ref. 109). After extracting our parameters from our combined DMC/DFT+U approach, we can plug those parameters into analytical models such as the one developed by Torelli and Olsen [160] to estimate the Curie temperature. We determined a maximum Tc of 43.56 K for CrI3 and 20.78 K for CrBr3, which is in excellent agreement with the measured values of 45 K [16] and 27 K, [180, 181] respectively. More details of this work can be found in Ref. 110.
@@ III.1.3 Monolayer MnO2
The workflow from Section III.1.2 is general and can be extended beyond CrX3 materials. In this work, we applied the same workflow to monolayer MnO2 [182]. Single layer MnO2 is a commercially available transition metal oxide semiconductor that has been synthesized and studied with DFT [183, 184, 185, 186, 187, 188, 35, 189]. Kan et al. [189] predicted the FM ordering to be more favorable than the AFM ordering with PBE+U (using a U correction of 3.9 eV obtained from previous literature [190]). was extracted from these calculations and a magnetic coupling Hamiltonian based on the Ising model was used to perform classical Monte Carlo simulations to obtain a Curie temperature of 140 K [189]. Although these PBE+U results are promising, there are aspects that can be revisited with more accurate methods such as DMC.
First, previous results and our own DFT+U benchmarking results show that the energy difference between the FM and AFM orientations is heavily dependent on the functional and U value [189, 182]. In fact, our DFT benchmarking shows an FM-AFM energy difference range of -55 meV to 18 meV, indicating that there is a discrepancy regardless of whether the FM or AFM orientation is more energetically favorable. Using the same workflow from Section III.1.2, we determined an optimal U value of 2.5 eV and a value of 1.2(5) meV with DMC (see Fig. 3). The starting geometry for our DMC calculations was obtained from the strongly constrained appropriately normed (SCAN) [66] meta-GGA functional (due to the fact that the SCAN computed lattice constant is in near-perfect agreement with experiment [183]). Using this optimal U value in subsequent PBE+U calculations, we find monolayer MnO2 to have out-of-plane magnetic anisotropy. From these magnetic constants, we estimated the maximum value of to be 28.8 K. In addition, we analyzed the spin density obtained from DMC for Mn and O and compared these with those obtained from PBE+U. Specifically, we find that DMC predicts that the spin density of O atoms is polarized antiparallel with respect to the Mn atoms, which is in agreement with PBE+U (U = 2.5, 3.5 eV), but the opposite of what is computed with PBE+U (U = 0, 1 eV). More details of this work can be found in Ref.182.

### v2

@@ II Theory
Although various real-space DMC codes have been developed over the past few decades and applied to 2D materials (e.g., CASINO, [177] PyQMC, [178] QWalk [179]), most of the work in this review article utilizes the QMCPACK code.[180, 181] The QMCPACK code has been routinely updated and maintained to track the latest algorithmic and hardware developments and utilize computational resources more efficiently (https://github.com/QMCPACK/qmcpack). The QMCPACK code is equipped with the Nexus [182] workflow automation software, which helps create and monitor DFT/DMC workflows that can significantly reduce the user time by tracking computational dependencies. More detailed information about these theoretical frameworks and codes can be found in Refs. 103, 63, 140, 141, 180, 181, 177, 178, 179, 183. For this review article, it is important to note that the error bars in each figure and the uncertainty noted in in concise notation (i.e., 5.3 0.1 can be expressed as 5.3(1)) refer to the standard error about the mean for each observable quantity calculated with QMC (e.g., energy, band gap, magnetic moment, lattice constant).
@@ III.1 Magnetic Properties
After the experimental synthesis of monolayer CrI3, which was measured to have a Curie temperature () of 45 K, [16] interest in identifying and physically understanding similar magnetic 2D materials has significantly increased. It has been experimentally reported that monolayer VSe2 is ferromagnetic (FM) on a vdW substrate [184] and that FM order exists in the bilayer limit for Cr2Ge2Te6.[185] Some other examples of low-dimensional magnetic systems include FePS3, [186] NiPS3, [187] MnPS3, [188] and Fe3GeTe2.[189] Several computational works have also predicted magnetism in a variety of 2D systems.[190, 191, 192, 193]
The Mermin-Wagner theorem [194] states that magnetic order in a 2D material cannot persist unless magnetic anisotropy is present. Moreover, it requires anisotropy that breaks continuous symmetries. Whether the anisotropy is perpendicular to the plane (easy axis) or parallel to the plane (easy plane) determines the type of transition temperature. For easy-axis anisotropy, the system can be described by the Heisenberg model with a finite Curie temperature . For easy-plane anisotropy, there is no explicit transition between the ordered and unordered states but instead a Kosterlitz-Thouless transition at finite temperature , where the system has quasi-long-range magnetic order below . To obtain accurate estimates for observables such as and , an accurate computation of the magnetic exchange and magnetic anisotropy energies is essential. Oftentimes, these results are extremely sensitive to the choice of density functional and Hubbard U parameter, resulting in inconclusive results with DFT. Additionally, these 2D magnetic materials can have a strong interdependence between structural parameters and magnetic properties, making it even more difficult to make accurate predictions. Due to this, using a method such as DMC to understand the electron correlation effects that drive magnetic ordering in 2D is necessary. These accurately computed ab initio magnetic parameters can then be used in conjunction with either classical Monte Carlo methods or other analytical models [195] to estimate observable quantities such as transition temperatures. Next, we discuss a few of our works on modeling magnetic 2D structures.
@@ III.1.1 Monolayer CrI3
One of the first monolayer magnets discovered,[16] monolayer CrI3 is a Mott insulator [196] that consists of chromium atoms octahedrally coordinated by six iodine atoms (see Fig. 1). Because of the negative exchange interactions among them, the spins on the Cr atoms arrange themselves perpendicular to the plane of the material,[197, 198] resulting in a monolayer ferromagnet with a critical temperature of 45 K.[16, 199] Upon stacking, CrI3 heterostructures oscillate between exhibiting ferromagnetism for odd numbers of layers and antiferromagnetism for even numbers of layers.[200] Moreover, upon the application of pressure that decreases the interlayer distance, bilayer CrI3 transitions from an antiferromagnet to a ferromagnet.[201, 202]
@@
These phenomena suggest that electron-electron and electron-phonon interactions play a key role in determining CrI3’s magnetic and structural properties. Yet, although significant research has gone into characterizing and modifying the macroscopic properties of this material, far less research has focused on the atomistic determinants of those properties in a manner that accurately accounts for many-body interactions. As a result, only approximate explanations for the source of this material’s magnetism had been put forth before the DMC research described below was conducted. In particular, based on a rudimentary analysis of oxidation states, it was originally believed that each iodine atom in CrI3 would have an oxidation state of 1 and therefore that the chromium should have an oxidation state of 3. Given that neutral Cr has a [Ar]4s13d5 electron configuration, this suggests that Cr3+ should have three valence electrons remaining, giving rise to a magnetic moment of 3 . More accurate yet still single-reference DFT calculations corroborated this picture—on average—but often differed significantly from one another in detail.[203, 204, 205, 197, 206] Previous PBE+U ( eV) and HSE calculations, for example, predicted Cr’s magnetic moment to be as large as 3.3 ,[204, 206] while PBE[206] and LDA+U [197] calculations predicted moments as small as 3.1 and 3 , respectively—a 10% difference among moments overall. Even larger differences could be observed among DFT predictions of the in-plane lattice parameter: PBE calculations predicted a CrI3 lattice parameter of 7.008 Å,[206] while LDA+U calculations predicted a lattice parameter of 6.686 Å,[197] a 5 % difference. These significant DFT discrepancies were further corroborated by our own DFT calculations, which demonstrated not only that the magnetic moments and lattice parameters predicted monotonically increased with the U employed in PBE+U calculations (suggesting that it would be difficult to determine a meaningful U value from DFT calculations alone) but also that the U that most closely matched experiments for one property significantly differed from those needed to match experimental predictions of other properties.[109] Thus, although these predictions provide general insights into CrI3’s magnetism, they lack the accuracy—and precision—needed to make definitive statements.
@@
To bring clarity to this picture, we thus performed DMC simulations on monolayer CrI3, one of the first performed on a magnetic monolayer material.[109] Given the known sensitivity of the electronic structure of 2D materials to strain[207, 208] and the lack of experimental results for the monolayer geometry, however, we first set out to determine a high-accuracy, many-body monolayer geometry (see Fig. 1). To do so, we leveraged a cutting-edge, many-body geometry optimization method: the surrogate Hessian line-search method.[209, 146, 174] In this approach, a Hessian generated using DFT is used to determine conjugate directions along which to locate the minimum-energy structure. Differing from standard DFT–based conjugate gradient minimization, the surrogate Hessian line-search method then computes DMC energies in parallel along each of the conjugate directions and subsequently refines the conjugate directions until the search converges to the DMC energy minimum.[209, 146, 174] This method is particularly advantageous because it makes the most efficient use of both DFT and DMC data: Instead of relying on more-expensive DMC energy gradients, it uses relatively cheap DFT gradients and then employs DMC energy calculations to refine the search, thus providing a substantially more efficient way of determining fully many-body geometries. Applying this approach to CrI3, we obtained a ground-state geometry with a lattice parameter of Å and a Cr-I bond distance of Å .[109] Interestingly, while we were conducting this research, some of the first scanning tunneling microscopy–based measurements of the monolayer geometry were performed, yielding a lattice constant of Å.[203] Such strong agreement (0.5% error) highlights the accuracy of our structure and the value of the surrogate Hessian line-search approach, despite CrI3’s (and many monolayer materials’) shallow potential energy surface (PES) around its minimum. This 0.5% error is quite remarkable considering that the DFT lattice constants had a 6 variation in answer (depending on the choice of DFT method).
@@
With this high-accuracy structure in hand, we then performed DMC calculations on the monolayer. To generate a high-quality trial wavefunction from which to construct our DMC nodal surface, we first performed PBE+U calculations. Unlike previous works, we were sure to determine our U in a variational fashion by determining the U in our PBE+U trial wavefunctions that minimized our resultant DMC energies (see Fig. 1). This resulted in an optimized U value of 2 eV.[109] Employing this U in our Slater-Jastrow trial wavefunctions with one- and two-body Jastrow terms, we obtained high-accuracy DMC charge and spin densities. We finally obtained DMC estimates of the Cr magnetic moment by integrating the spin density from the center of the Cr atom to the distance from the nucleus at which the spin density changed sign (the zero-recrossing radius of the sign of the spin density). This yielded a DMC magnetic moment of 3.62 for Cr atom, a moment substantially larger than that previously predicted using single-reference techniques. Importantly, PBE+U 2 eV calculations performed on the same optimized geometries yielded a similarly large Cr moment of 3.57 , demonstrating the strong coupling between the moment and the geometry,[208, 207] as also observed in previous DFT simulations and our subsequent spin-phonon and spin-lattice coupling calculations.[109] Although these moments seem unexpectedly large, they are counterbalanced by negative moments of 0.145 on the iodine atoms, leading to the anticipated averaged moment of 3 over the entire unit cell. These simulations have therefore shed a clarifying light on the atomic-scale origin of CrI3’s magnetism and produced a valuable workflow that can be readily extended to a wide range of other 2D materials.[210] Additional details of this work can be found in Ref. 109.
@@ III.1.2 Monolayer CrX3
To expand upon this work, we created a workflow that combines DFT+U and DMC methods to calculate magnetic properties of 2D CrX3 (X = I, Br, Cl, F) with improved accuracy.[110] These structures were chosen as a case study because they have been extensively investigated with DFT, [211] have been experimentally synthesized, and have a finite transition temperature.[16, 211] We can map our ab initio quantities to a 2D model spin Hamiltonian to obtain properties such as .[195, 197] The model Hamiltonian contains easy-axis single-ion anisotropy (), Heisenberg isotropic exchange (), and anisotropic exchange (),
@@
The sum over runs over the lattice of Cr atoms (magnetic atoms), and runs over ’s nearest-neighbor Cr atoms; only nearest neighbors are needed because the magnetic moments are strongly localized on the Cr atoms. prefers off-plane easy axis, prefers FM interactions, and results in fully isotropic exchange. Note that in the following sections, refers to the Heisenberg exchange and not the J in the DFT+U+J formalism.[212] Two magnetic orientations can be constructed from the CrX3 unit cell: the FM (two spin-up Cr atoms) and antiferromagnetic (AFM) (one spin-up Cr atom and one spin-down Cr atom) orientations. , , and can be extracted from noncollinear (spin-orbit) DFT calculations. Specifically, the energy differences between configurations that have an easy-axis rotation of 90∘ and nonrotated configurations for FM and AFM magnetic orientations can be computed.
@@
These results can be systematically improved with DMC. First, the nodal surface of the trial wavefunction can be optimized by changing the U value (variationally determining which U value yields the lowest energy). Second, by performing DMC for the FM and AFM configurations, can be estimated. In our case, these DMC energies are collinear (without SOC). Although SOC has been implemented in DMC, the energy differences needed to compute magnetic anisotropy parameters are smaller in magnitude than the DMC uncertainty. For this reason, we used the optimal U value determined from the nodal surface optimization (in our case, it was eV for CrI3 and CrBr3, similar to Ref. 109) to perform DFT+U calculations to obtain and . We performed this DMC variational optimization of U for both the FM and AFM configurations of 2D CrI3 and CrBr3, where both magnetic configurations yielded an optimal U value of 2 eV. The starting geometry for our DMC calculations was obtained from the vdW-DF-optB88 [79] functional (vdW-DF-optB88 produces a geometry for CrI3 in near-perfect agreement with experiment [203] and Ref. 109). After extracting our parameters from our combined DMC/DFT+U approach, we can plug those parameters into analytical models such as the one developed by Torelli and Olsen [195] to estimate the Curie temperature. We determined a maximum of 43.56 K for CrI3 and 20.78 K for CrBr3, which is in excellent agreement with the measured values of 45 K [16] and 27 K, [213, 214] respectively. More details of this work can be found in Ref. 110.
@@ III.1.3 Monolayer MnO2
The workflow from Section III.1.2 is general and can be extended beyond CrX3 materials. In this work, we applied the same workflow to monolayer MnO2.[175] Single-layer MnO2 is a commercially available transition metal oxide semiconductor that has been synthesized and studied with DFT.[215, 216, 217, 218, 219, 220, 35, 221] Kan et al. [221] predicted the FM ordering to be more favorable than the AFM ordering with PBE+U (using a U correction of 3.9 eV obtained from previous literature [222]). was extracted from these calculations, and a magnetic coupling Hamiltonian based on the Ising model was used to perform classical Monte Carlo simulations to obtain a Curie temperature of 140 K.[221] Although these PBE+U results are promising, there are aspects that can be revisited with more accurate methods such as DMC.
First, previous results and our own DFT+U benchmarking results show that the energy difference between the FM and AFM orientations is heavily dependent on the functional and U value.[221, 175] In fact, our DFT benchmarking shows an FM-AFM energy difference range of 55 meV to 18 meV, indicating that there is a discrepancy regardless of whether the FM or AFM orientation is more energetically favorable. Using the same workflow from Section III.1.2, we determined an optimal U value of 2.5 eV and a value of 1.2(5) meV with DMC (see Fig. 2). The starting geometry for our DMC calculations was obtained from the strongly constrained appropriately normed (SCAN) [66] meta-GGA functional (because the SCAN-computed lattice constant is in near-perfect agreement with experiment [215]). Using this optimal U value in subsequent PBE+U calculations, we find monolayer MnO2 to have out-of-plane magnetic anisotropy. From these magnetic constants, we estimated the maximum value of to be 28.8 K. In addition, we analyzed the spin density obtained from DMC for Mn and O and compared these with those obtained from PBE+U. Specifically, we found that DMC predicts that the spin density of O atoms is polarized antiparallel with respect to the Mn atoms, which is in agreement with our PBE+U results for eV and eV (the O spin density is polarized parallel with respect to Mn for eV and eV). More details of this work can be found in Ref. 175.